OpenStudy (anonymous):

a piece of copper has a volume of 660L. how many atoms does this sample contain?

OpenStudy (abb0t):

Convert to mol using the relations 22.7 L/mol assuming STP

OpenStudy (abb0t):

then convert to atoms using avogadros number, \(6..022 \times 10^{23} \frac{atoms}{mol}\)

OpenStudy (anonymous):

got it thank you

OpenStudy (anonymous):

NO!!! 22.7 mol/L only holds at STP for *gases*. Copper is not going to be a gas at STP.

OpenStudy (anonymous):

Use the density of copper \(\rho=8.96\text{ g cm}^{-3}=8.96\text{ g/mL}=8960\text{ g/L}\). So we can determine the mass of our sample easily$$660\text{ L}\times8960\text{ g/L}=5900000\text{ g} $$Given that we know the molar mass of copper to be \(63.55\text{ g/mol}\) we can determine our number of moles too:$$5913600\text{ g}\div63.55\text{ g/mol}=93000\text{ mol}$$To count the number of atoms, recall there are \(N_A\) atoms per mole:$$93000\text{ mol}\times6.022\times10^{23}\text{ atoms/mol}=5.6\times10^{28}\text{ atoms}$$

OpenStudy (abb0t):

Ah, you're right. Thank you.

OpenStudy (abb0t):

My mistake.